OpenStudy (anonymous):

Solve the System of Equations by using the Elimination Method:

OpenStudy (anonymous):

\[x^2+y^2=3 \] \[4x^2+y=0 \]

OpenStudy (anonymous):

so that will get me \[-4y^2 +y=-12]\ ?

OpenStudy (anonymous):

\[4y^2-y=12 \] so then it do I divide the 4 and is it \[y^2-y=3\] then I subtract the 3 and get\[y^2-y-3=0\] that way I can factor it?

OpenStudy (anonymous):

oh okay then what do I do after that?

OpenStudy (anonymous):

oh okay